# Fully Sparse 3D Occupancy Prediction

### v3

@@ 5 Experiments
We evaluate our model on the Occ3D-nus [48] dataset. Occ3D-nus is based on the nuScenes [3] dataset, which consists of large-scale multimodal data collected from 6 surround-view cameras, 1 LiDAR sensor and 5 radar sensors. The dataset has 1000 videos and is split into 700/150/150 videos for training/validation/testing. Each video has roughly 20s duration and the key samples are annotated at intervals of 0.5s.
@@
| Method | Backbone | Input Size | Epochs | RayIoU | RayIoU1m | RayIoU2m | RayIoU4m | FPS |
| BEVFormer [26] | R101 | 1600 900 | 24 | 32.4 | 26.1 | 32.9 | 38.0 | 3.0 |
| RenderOcc [39] | Swin-B | 1408 512 | 12 | 19.5 | 13.4 | 19.6 | 25.5 | - |
| BEVDet-Occ [13] | R50 | 704 256 | 90 | 29.6 | 23.6 | 30.0 | 35.1 | 2.6 |
| BEVDet-Occ-Long (8f) | R50 | 704 384 | 90 | 32.6 | 26.6 | 33.1 | 38.2 | 0.8 |
| FB-Occ (16f) [27] | R50 | 704 256 | 90 | 33.5 | 26.7 | 34.1 | 39.7 | 10.3 |
| SparseOcc (8f) | R50 | 704 256 | 24 | 34.0 | 28.0 | 34.7 | 39.4 | 17.3 |
| SparseOcc (16f) | R50 | 704 256 | 24 | 35.1 | 29.1 | 35.8 | 40.3 | 12.5 |
@@ 5.1 Implementation Details
During training, we use the AdamW [35] optimizer with a global batch size of 8. The initial learning rate is set to and is decayed with cosine annealing policy. For all experiments, we train our models for 24 epochs. FPS is measured on a Tesla A100 GPU with the PyTorch fp32 backend (single batch size).
@@ 5.2 Main Results
In Tab. 1 and Fig. 1 (c), we compare SparseOcc with previous state-of-the-art methods on the validation split of Occ3D-nus. Despite under a weaker setting (ResNet-50 [12], 8 history frames, and input image resolution of 704 256), SparseOcc significantly outperforms previous methods including FB-Occ, the winner of CVPR 2023 occupancy challenge, with many complicated designs including forward-backward view transformation, depth net, joint depth and semantic pre-training, and so on. SparseOcc achieves better results (+1.6 RayIoU) while being faster and simpler than FB-Occ, which demonstrates the superiority of our solution.
@@ 5.3 Ablations
In Tab. 3, we first ablate the effectiveness of mask transformer. The first row is a per-voxel baseline which directly predicts semantics from the sparse voxel decoder using a stack of MLPs. Introducing mask transformer with vanilla dense cross attention (as it is the common practice in MaskFormer and Mask3D) gives a performance boost of 1.7 RayIoU, but inevitably slows down the inference speed. Therefore, to speed up the dense cross-attention pipeline, we adopt a sparse sampling mechanism which brings a 50% reduction in inference time. Then, by further sampling sparse 3D points via predicted mask guidance, we finally achieve 29.2 RayIoU with 24 FPS.
@@
In this study, we delve deeper into the impact of voxel sparsity on final performance. To investigate this, we systematically ablate the value of in Fig. 8 (a). Starting from a modest value of 16000, we observe that the optimal performance occurs when is set to 32000, which is only 5% of the total number of dense voxels (20020016 = 640000). Surprisingly, further increasing does not yield any performance improvements; instead, it introduces noise. Thus, our findings suggest that a 5% sparsity level is sufficient, and additional sparsity would be counterproductive.
@@
In Fig. 8 (b), we validate the effectiveness of temporal fusion. We can see that the temporal modeling of SparseOcc is very effective, with performance steadily increasing as the number of frames increases. The performance peaks at 12 frames and then saturates.
5.4 Extensive Experiments
Panoptic occupancy.
We then show that SparseOcc can be easily extended for panoptic occupancy predictions, a task derived from panoptic segmentation that segments images to not only semantically meaningful regions but also to detect and distinguish individual instances. Compared to panoptic segmentation, panoptic occupancy prediction requires the model to be geometry awareness to construct the 3D scene for segmentation. By additionally introducing instance queries to the mask transformer, we seamlessly achieve the first panoptic occupancy prediction framework using camera inputs. In Fig. 9, we visualize the panoptic occupancy results of SparseOcc. As illustrated, semantic regions and individual objects as well as their occupancy grids are well predicted. For more details, please refer to the appendix.
Enhancing sparsity by removing the road surface.
The majority of non-free occupancy data pertains to background geometry. In practice, the drivable surface occupancy can be effectively substituted with High-Definition Map (HD Map) or online mapping techniques [5, 28, 50, 24]. This replacement not only streamlines the sparsity but also enriches the semantic and structural understanding of roads. We construct experiments to investigate the effect of removing road surface in SparseOcc. The details can be found in the appendix.
6 Limitations
Accumulative errors.
7 Conclusion
In this paper, we have proposed a fully sparse occupancy network, named SparseOcc, which neither relies on dense 3D feature, nor has sparse-to-dense and global attention operations. We also created RayIoU, a ray-level metric for occupancy evaluation, eliminating the inconsistency flaws of previous metric. Experiments show that SparseOcc achieves the state-of-the-art performance on Occ3D-nuScenes for both speed and accuracy. We hope this exciting result will attract more attention to the fully sparse 3D occupancy paradigm.
Appendix A More Results
| Method | Backbone | Input Size | Epochs | Vis. Mask | RayIoU | mIoU | FPS |
| RenderOcc [39] | Swin-B | 1408 512 | 12 | 19.5 | 24.4 | - | |
| BEVDet-Occ [13] | R50 | 704 256 | 90 | 29.6 | 36.1 | 2.6 | |
| BEVDet-Occ-Long (8f) | R50 | 704 384 | 90 | 32.6 | 39.3 | 0.8 | |
| BEVFormer [26] | R101 | 1600 900 | 24 | 32.4 | 39.3 | 3.0 | |
| FB-Occ (16f) [27] | R50 | 704 256 | 90 | 33.5 | 39.1 | 10.3 | |
| BEVFormer [26] | R101 | 1600 900 | 24 | - | 33.7 | 23.7 | 3.0 |
| FB-Occ (16f) [27] | R50 | 704 256 | 90 | - | 35.6 | 27.9 | 10.3 |
| SparseOcc (16f) | R50 | 704 256 | 24 | - | 35.1 | 30.6 | 12.5 |
| SparseOcc (16f) | R50 | 704 256 | 48 | - | 36.1 | 30.9 | 12.5 |
In this section, we present a more detailed comparison in Tab. 4, which includes performance metrics based on the traditional voxel-level mIoU. Our model, SparseOcc, does not utilize the visible mask during training. To ensure a fair comparison, we trained a variant of FB-Occ [27] that also does not use the visible mask during training. Despite operating under a weaker setting (48 vs. 90 epochs), our SparseOcc still surpasses FB-Occ in both accuracy and speed. Besides, FB-Occ employs many complicated designs, tricks and data augmentations, while our SparseOcc is more concise and elegant.
Interestingly, we observed a peculiar phenomenon. Under the traditional voxel-level mIoU metric, methods can significantly benefit from disregarding the non-visible voxels during training. These non-visible voxels are indicated by a binary visible mask provided by the Occ3D-nuScenes dataset. However, this strategy actually impairs performance under our new RayIoU metric. For instance, we train two variants of BEVFormer [26]: one uses the visible mask during training, and the other does not. The former scores 15 points higher than the latter on the voxel-based mIoU, but it scores 1 point lower on RayIoU. This phenomenon is also observed on FB-Occ.
@@ Appendix A More Results
| BEVFormer | 33.7 | 5.0 | 42.2 | 18.2 | 55.2 | 57.1 | 22.7 | 21.3 | 31.0 | 27.1 | 30.7 | 49.4 | 58.4 | 30.4 | 29.4 | 31.7 | 36.3 | 26.5 |
| BEVFormer | 32.4 | 6.4 | 44.8 | 24.0 | 55.2 | 56.7 | 21.0 | 29.8 | 33.5 | 26.8 | 27.9 | 49.5 | 45.8 | 18.7 | 22.4 | 18.5 | 39.1 | 29.8 |
| FB-Occ | 35.6 | 10.5 | 44.8 | 25.6 | 55.6 | 51.7 | 22.6 | 27.2 | 34.3 | 30.3 | 23.7 | 44.1 | 65.5 | 33.3 | 31.4 | 32.5 | 39.6 | 33.3 |
| FB-Occ | 33.5 | 5.0 | 44.9 | 26.2 | 59.7 | 55.1 | 27.9 | 29.1 | 34.3 | 29.6 | 29.1 | 50.5 | 44.4 | 22.4 | 21.5 | 19.5 | 39.3 | 31.1 |
@@
To explore this phenomenon, we present the per-class RayIoU in Tab. 5. The table reveals that using the visible mask during training enhances performance for most foreground classes such as bus, bicycle, and truck. However, it negatively impacts background classes like drivable surface, terrain, and sidewalk.
@@
This observation prompts a further question: why does performance deteriorate for background classes? To address this, we offer a visual comparison of depth errors and height maps of the predicted drivable surface from FB-Occ, both with and without the use of visible mask during training, in Fig. 10. The figure illustrates that training with visible mask results in a thicker and higher ground representation, leading to substantial depth errors in distant areas. Conversely, models trained without the visible mask predict depth with greater accuracy.
@@
From these observations, we derive some valuable insights: ignoring non-visible voxels during training benefits foreground classes by resolving the issue of ambiguous labeling of unscanned voxels. However, it also compromises the accuracy of depth estimation, as models tend to predict a thicker and closer surface. We hope that our findings will inform and benefit future research.
Appendix B Panoptic Occupancy Prediction
Thanks to the mask transformer, our SparseOcc can produce panoptic occupancy prediction by simply replacing the semantic queries with instance queries.
Ground Truth Preparation.
@@ Appendix B Panoptic Occupancy Prediction
To evaluate our method, we utilize the ground-truth object bounding boxes from the 3D detection task to generate the panoptic occupancy ground truth. First, we define eight instance categories (including car, truck, construction vehicle, bus, trailer, motorcycle, bicycle, pedestrian) and ten staff categories (including terrain, manmade, vegetation, etc). Next, we identify each instance segment by grouping the voxels inside the box based on an existing semantic occupancy benchmark Occ3D-nuScenes [48].
However, we observe that using the original size of the box for grouping may cause some boundary voxels being missed due to the compactness of the box. Enlarging the box (such as 1.2x) leads to excessive overlap between boxes. To address these issues, we designed a two-stage grouping scheme. In the first stage, we use the original size of the box for grouping. Then, for voxels that have not been assigned to a specific instance, we select the closest box and assign it. This scheme effectively resolves the problems of boundary omission and box overlap.
Evaluation Metrics.
@@
We design RayPQ based on the well-known panoptic quality (PQ) [20] metric, which is defined as the multiplication of segmentation quality (SQ) and recognition quality (RQ):
@@
| Method | Backbone | Input Size | Epochs | RayPQ | RayPQ1m | RayPQ2m | RayPQ4m |
@@
Results.
In Tab. 6, we report the performance of SparseOcc on panoptic occupancy benchmark. Similar to RayIoU, we calculate RayPQ under three distance thresholds: 1, 2 and 4 meters. SparseOcc achieves an averaged RayPQ of 14.1. The visualizations are presented in the main paper (Fig. 9).
Appendix C Enhancing Sparsity
As mentioned in the main paper, the majority of non-free occupancy data pertains to the background geometry, such as the road surface. In practice, the occupancy of road surface can be effectively substituted with High-Definition Map (HD Map) or online mapping techniques [5, 28, 50, 24]. Thus, the sparsity of the scene can be further enhanced by removing certain background categories, leading to faster inference speed with negligible performance loss. This is also an advantage of SparseOcc compared to the dense counterparts, because the dense methods will not speed up as the sparsity of the scene increases.
| Method | Backbone | Input Size | Epochs | Top- | RayIoU* | FPS |
| SparseOcc | R50 | 704 256 | 24 | 32000 | 30.1 | 24.0 |
| SparseOcc | R50 | 704 256 | 24 | 24000 | 29.8 | 24.8 |
| SparseOcc | R50 | 704 256 | 24 | 16000 | 28.8 | 26.0 |
| SparseOcc | R50 | 704 256 | 24 | 32000 | 30.1 | 24.0 |
| SparseOcc | R50 | 704 256 | 24 | 24000 | 30.0 | 24.8 |
| SparseOcc | R50 | 704 256 | 24 | 16000 | 29.4 | 26.0 |
Settings.
We train a variant of the model that the voxels belonging to the drivable surface and terrian in the ground truth are treated as free during training (denoted by in Tab. 7). For fair evaluation, all models are evaluated on the categories that are not ignored.
Results.
As shown in Tab. 7, the performance of baseline (modeling all categories) drops notably as the top- decreases. This is reasonable as the number of voxels is not enough to express the entire scene. In contrast, if we ignore certain background categories, the performance loss is negligible (only 0.7 RayIoU) even when top- is reduced by half. This means the inference speed of SparseOcc can be further improved by enhancing sparsity, while for the dense counterparts it is not possible.
Appendix D Visualization of 3D Reconstruction
In Fig. 11, we visualize the reconstructed 3D geometry from sparse voxel decoder. SparseOcc can reconstruct fine-grained details from camera-only inputs.

### v4

@@ 5 Experiments
We evaluate our model on the Occ3D-nuScenes [48] dataset. Occ3D-nuScenes is based on the nuScenes [3] dataset, which consists of large-scale multimodal data collected from 6 surround-view cameras, 1 lidar and 5 radars. The dataset has 1000 videos in total and is split into 700/150/150 videos for training/validation/testing. Each video has roughly 20s duration and the key samples are annotated every 0.5s.
@@
| Method | Backbone | Input Size | Epoch | RayIoU | RayIoU1m, 2m, 4m | mIoU | FPS | ||
| BEVFormer (4f) [26] | R101 | 1600900 | 24 | 32.4 | 26.1 | 32.9 | 38.0 | 39.2 | 3.0 |
| RenderOcc [38] | Swin-B | 1408512 | 12 | 19.5 | 13.4 | 19.6 | 25.5 | 24.4 | - |
| SimpleOcc [11] | R101 | 672336 | 12 | 22.5 | 17.0 | 22.7 | 27.9 | 31.8 | 9.7 |
| BEVDet-Occ (2f) [14] | R50 | 704256 | 90 | 29.6 | 23.6 | 30.0 | 35.1 | 36.1 | 2.6 |
| BEVDet-Occ-Long (8f) | R50 | 704384 | 90 | 32.6 | 26.6 | 33.1 | 38.2 | 39.3 | 0.8 |
| FB-Occ (16f) [27] | R50 | 704256 | 90 | 33.5 | 26.7 | 34.1 | 39.7 | 39.1 | 10.3 |
| SparseOcc (8f) | R50 | 704256 | 24 | 34.0 | 28.0 | 34.7 | 39.4 | 30.1 | 17.3 |
| SparseOcc (16f) | R50 | 704256 | 24 | 35.1 | 29.1 | 35.8 | 40.3 | 30.6 | 12.5 |
| SparseOcc (16f) | R50 | 704256 | 48 | 36.1 | 30.2 | 36.8 | 41.2 | 30.9 | 12.5 |
@@ 5.1 Implementation Details
During training, we use the AdamW [34] optimizer with a global batch size of 8. The initial learning rate is set to and is decayed with cosine annealing policy. For all experiments, we train our models for 24 epochs. FPS is measured on a Tesla A100 GPU with the PyTorch fp32 backend.
@@ 5.2 Main Results
In Tab. 1 and Fig. 1 (b), we compare SparseOcc with previous state-of-the-art methods on the validation split of Occ3D-nuScenes. Despite under a weaker setting (ResNet-50 [13], 8 history frames, and input image resolution of 704 256), SparseOcc significantly outperforms previous methods including FB-Occ, the winner of CVPR 2023 occupancy challenge, with many complicated designs including forward-backward view transformation, depth net, joint depth and semantic pre-training, and so on. SparseOcc achieves better results (+1.6 RayIoU) while being much faster and simpler than FB-Occ, which demonstrates the superiority of our solution.
@@ 5.3 Ablations
In Tab. 3, we ablate the effectiveness of the mask transformer. The first row is a simple per-voxel baseline which directly predicts semantics from the sparse voxel decoder using a stack of MLPs. Introducing mask transformer with vanilla cross attention (as it is the common practice in MaskFormer and Mask3D) gives a performance boost of 1.7 RayIoU, but inevitably slows down the inference speed as it attends to all locations in an image. Therefore, to speed up the dense cross-attention pipeline, we adopt a sparse sampling mechanism which brings a 50% reduction in inference time. By further introducing the predicted masks to guide the generation of sampling points, we finally achieve 29.2 RayIoU with 24 FPS.
@@
In this study, we delve deeper into the impact of voxel sparsity on final performance. To investigate this, we systematically ablate the value of in Fig. 8 (a). Starting from a modest value of 16k, we observe that the optimal performance occurs when is set to 32k 48k, which is only 5% 7.5% of the total number of dense voxels (20020016 = 640000). Surprisingly, further increasing does not yield any performance improvements; instead, it introduces noise. Thus, our findings suggest that a 5% sparsity level is sufficient. Keep increasing the density will reduce both accuracy and speed.
Pruning by top- is simple and effective, but it is related to specific dataset. In real world, we can substitute top- with a thresholding method. Voxels scoring less than a given threshold (e.g., 0.7) will be pruned. Thresholding achieves similar performance to top- (see Fig. 8 (b)), and has the ability to generalize to different scenes.
@@
In Fig. 8 (c), we validate the effectiveness of temporal fusion. We can see that the temporal modeling of SparseOcc is very effective, with performance steadily increasing as the number of frames increases. The performance peaks at 12 frames and then saturates. However, the inference speed drops rapidly as the sampling points need to interact with every frame.
5.4 More Studies
| Per-class RayIoU | |||||||||||||||||||
@@ 5.4 More Studies
| BEVFormer | 23.7 | 33.7 | 5.0 | 42.2 | 18.2 | 55.2 | 57.1 | 22.7 | 21.3 | 31.0 | 27.1 | 30.7 | 49.4 | 58.4 | 30.4 | 29.4 | 31.7 | 36.3 | 26.5 |
| BEVFormer | 39.2 | 32.4 | 6.4 | 44.8 | 24.0 | 55.2 | 56.7 | 21.0 | 29.8 | 33.5 | 26.8 | 27.9 | 49.5 | 45.8 | 18.7 | 22.4 | 18.5 | 39.1 | 29.8 |
| FB-Occ | 27.9 | 35.6 | 10.5 | 44.8 | 25.6 | 55.6 | 51.7 | 22.6 | 27.2 | 34.3 | 30.3 | 23.7 | 44.1 | 65.5 | 33.3 | 31.4 | 32.5 | 39.6 | 33.3 |
| FB-Occ | 39.1 | 33.5 | 5.0 | 44.9 | 26.2 | 59.7 | 55.1 | 27.9 | 29.1 | 34.3 | 29.6 | 29.1 | 50.5 | 44.4 | 22.4 | 21.5 | 19.5 | 39.3 | 31.1 |
The effect of training with visible masks.
Interestingly, we observed a peculiar phenomenon. Under the traditional voxel-level mIoU metric, dense methods can significantly benefit from disregarding the non-visible voxels during training. These non-visible voxels are indicated by a binary visible mask provided by the Occ3D-nuScenes dataset. However, we find that this strategy actually impairs performance under our new RayIoU metric. For instance, we train two variants of BEVFormer: one uses the visible mask during training, and the other does not. As shown in Tab. 4, the former scores 15 points higher than the latter on the voxel-based mIoU, but it scores 1 point lower on RayIoU. This phenomenon is also observed on FB-Occ.
@@
To further explore this, we present the per-class RayIoU in Tab. 4. The table reveals that training with the visible mask enhances performance for most foreground classes such as bus, bicycle, and truck. However, it negatively impacts background classes like drivable surface and terrain.
@@
This observation raises a further question: Why does the performance of the background category degrade? To address this, we offer a visual comparison of the depth errors and height maps of the predicted drivable surface from FB-Occ in Fig. 9, both with and without the use of visible mask during training. The figure illustrates that training with visible masks results in a thicker and higher ground prediction, leading to substantial depth errors in distant areas. Conversely, models trained without the visible mask predict depth with greater accuracy.
@@
From these observations, we derive some valuable insights: ignoring non-visible voxels during training benefits foreground classes by resolving the issue of ambiguous labeling of unscanned voxels. However, it also compromises the accuracy of depth estimation, as models tend to predict a thicker and closer surface. We hope that our findings will benefit future research.
Panoptic occupancy.
We then show that SparseOcc can be easily extended for panoptic occupancy prediction, a task derived from panoptic segmentation that segments images to not only semantically meaningful regions but also to detect and distinguish individual instances. Compared to panoptic segmentation, panoptic occupancy prediction requires the model to have geometric awareness in order to construct the 3D scene for segmentation. By additionally introducing instance queries to the mask transformer, we seamlessly achieve the first fully sparse panoptic occupancy prediction framework using camera-only inputs.
@@
Firstly, we utilize the ground-truth bounding boxes from the 3D object detection task to generate the panoptic occupancy ground truth. Specifically, we define eight instance categories (including car, truck, construction vehicle, bus, trailer, motorcycle, bicycle, pedestrian) and ten staff categories (including terrain, manmade, vegetation, etc). Each instance segment is identified by grouping the voxels inside the bounding box based on an existing semantic occupancy benchmark, such as Occ3D-nuScenes.
@@
We then design RayPQ based on the well-known panoptic quality (PQ) [21] metric, which is defined as the multiplication of segmentation quality (SQ) and recognition quality (RQ):
@@
| Method | Backbone | Input Size | Epoch | RayPQ | RayPQ1m | RayPQ2m | RayPQ4m |
@@
5.5 Limitations
Accumulative errors.
6 Conclusion
In this paper, we proposed a fully sparse occupancy network, named SparseOcc, which neither relies on dense 3D feature, nor has sparse-to-dense and global attention operations. We also created RayIoU, a ray-level metric for occupancy evaluation, eliminating the inconsistency flaws of previous metric. Experiments show that SparseOcc achieves the state-of-the-art performance on the Occ3D-nuScenes dataset for both speed and accuracy. We hope this exciting result will attract more attention to the fully sparse 3D occupancy paradigm.
Acknowledgements
We thank the anonymous reviewers for their suggestions that make this work better. This work is supported by the National Key RD Program of China (No. 2022ZD0160900), the National Natural Science Foundation of China (No. 62076119, No. 61921006), the Fundamental Research Funds for the Central Universities (No. 020214380119), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.